# Convergence Rate of Particle System for Second-order PDEs On Wasserstein Space

### v1

@@ 1. Introduction
$\begin{cases}-\partial_{t}v^{N}(t,\boldsymbol{x})=\frac{1}{N}\sum_{i=1}^{N}H(x^{i},ND_{x^{i}}v^{N}(t,\boldsymbol{x}),\mu^{\boldsymbol{x}})+\sum_{i=1}^{N}\Delta_{x^{i}}v^{N}(t,\boldsymbol{x})\\ \quad\quad\quad\quad\quad\quad\ \ +a\sum_{i,j=1}^{N}\operatorname{tr}(D_{x^{i}x^{j}}v^{N})(t,\boldsymbol{x}),\\ \quad\ v^{N}(T,\boldsymbol{x})=G(\mu^{\boldsymbol{x}})\end{cases}$ (1.2)
@@
$\begin{cases}-\partial_{t}V=\int_{\mathbb{T}^{d}}H^{e}(x,z,D_{\mu}V(t,z,\mu,x),\mu)\,\mu(dx)+\int_{\mathbb{T}^{d}}\operatorname{tr}(D_{x\mu}^{2}V)(t,z,\mu,x)\,\mu(dx)\\ \quad\quad\quad\quad\quad+a\Delta_{z}V(t,z,\mu),\\ \quad V(T,z,\mu)=G^{e}(z,\mu),\end{cases}$ (1.3)
@@
is a viscosity supersolution to (1.3) up to some error term $E(\epsilon,N)$, and therefore by comparison
$V^{N}(t,z,\boldsymbol{x})+E(\epsilon,N)\geq\overline{V}^{N,\epsilon}(t,z,\mu^{\boldsymbol{x}})+E(\epsilon,N)\geq V(t,z,\mu^{\boldsymbol{x}}),\quad(t,z,\boldsymbol{x})\in[0,T]\times\mathbb{T}^{d+dN}.$
@@
Then choosing $\epsilon$ properly as a function of $N$, one obtains that $E(\epsilon,N)\approx C\alpha(N)$ where $C$ is a positive constant independent of $N$ and
$\alpha(N)=\begin{cases}N^{-1/2},\quad&\text{if $d\in(0,2)$},\\ N^{-1/2}\log(N),&\text{if $d=2$},\\ N^{-1/d},&\text{if $d>2$}.\end{cases}$
@@
    $G$ is Lipschitz with respect to $C^{-k_{*}}$;
@@
    For each $R>0$, there is a constant $C_{R}$ such that for each $(x,p)\in\mathbb{T}^{d}\times B_{R}$
@@
Under Assumption 1.1, $v^{N}$ converges to $v$ with the rate $\alpha(N)$
$\sup_{(t,\boldsymbol{x})\in[0,T]\times\mathbb{T}^{dN}}\left|v^{N}(t,\boldsymbol{x})-v(t,\mu^{\boldsymbol{x}})\right|\leq C\alpha(N),$
@@
Our result provides the first convergence rate for particle approximations of partially second-order PDEs on Wasserstein space. Our argument is purely analytic and doesn’t rely on the semi-concavity/convexity of $H$. We would like to mention that the argument of our result makes use of the regularity of $v^{N}$ established in [19] where only the convergence to 00 of $\sup_{(t,\boldsymbol{x})\in[0,T]\times\mathbb{T}^{dN}}\left|v^{N}(t,\boldsymbol{x})-v(t,\mu^{\boldsymbol{x}})\right|$ is provided and not its convergence rate. Moreover, using the same method of the infinite dimensional inf/sup-convolution and the comparison result from [4], one should also be able to show the convergence rate for the same type of equations on ${\mathbb{R}}^{d}$. But to illustrate the main idea of the argument, we choose to work on $\mathbb{T}^{d}$ to avoid the technical issue of non-compactness.
@@ 1.1. Related Literature
Convergence of particle system in mean field control problems were studied in [35, 34, 2] based on viscosity theory, while [26] provided a probabilistic argument. The convergence rate for first-order PDEs on Wasserstein space was obtained in [11, 18, 9]. Assuming the existence of smooth solution to mean-field PDEs, [24] got the optimal convergence rate by a verification argument.
@@ 1.2. Notations
Throughout the paper, we adopt Lions differentiability for functions defined on ${\mathcal{P}}(\mathbb{T}^{d})$; see e.g. [12, Chapter 5].
@@ 2.1. Definition of viscosity solution
Similarly, we define viscosity supersolution. A continuous function $v:[0,T]\times{\mathcal{P}}(\mathbb{T}^{d})\to{\mathbb{R}}$ is called a viscosity solution if its extension $V$ via (2.1) is a viscosity subsolution and supersolution to (1.3) at the same time.
@@ 2.2. Preliminary results
As the HJB equation (1.2) of particle system is uniformly elliptic, one can show the regularity of solution. The following result is from [19, Lemma 3.1, Theorem 3.2].
@@
$\sup_{(t,\mu)\in{[0,T]\times\mathbb{T}^{d}}}\lVert\frac{\delta\hat{v}^{N}}{\delta\mu}(t,\mu,\cdot)\rVert_{\mathbb{H}^{k_{*}}}\leq\sum_{k=1,\dotso,k_{*}}\sum_{i=1}^{N}\lVert D_{x^{i}}^{k}v^{N}\rVert_{\infty}\leq C.$
@@
By Arzelà–Ascoli theorem, Lemma 2.2 implies there exists a subsequence of $\hat{v}^{N}$ uniformly converging to a function $v$, which is also $\rho_{*}$-Lipschitz. Moreover according to Proposition 2.1, $v$ is also a limiting point of $v^{N}$, and hence is a viscosity solution to (1.1) by a standard argument. Finally, the uniqueness of viscosity solution has been proved in [19] based on the Ishii’s Lemma in [4].
@@
For (1.1), any upper semi-continuous subsolution is smaller than lower semi-continuous supersolution. Then there is a unique viscosity solution $v$ to (1.1) on $\mathcal{P}(\mathbb{T}^{d})$.
@@
Thanks to Sobolev embedding theorem [7, Corollary 9.13], as $k_{*}=\lfloor d/2\rfloor+2$ any $f\in\mathbb{H}^{k_{*}}(\mathbb{T}^{d})$ is Lipschitz with coefficient proportional to $\lVert f\rVert_{\mathbb{H}^{k_{*}}}$. Therefore, it can be easily seen that
@@
$V^{N}(t,z,\boldsymbol{x})\geq V(t,z,\mu^{\boldsymbol{x}})-C\alpha(N),$ (3.1)
@@
$V^{N}(t,z,\boldsymbol{x})\leq V(t,z,\mu^{\boldsymbol{x}})+C\alpha(N)$
@@
$-\partial_{t}\Phi(t_{0},z_{0},\mu_{0})+E(N,\epsilon)\geq$ $\int_{\mathbb{T}^{d}}H^{e}(x,z_{0},D_{\mu}\Phi(t_{0},z_{0},\mu_{0},x),\mu_{0})\,\mu_{0}(dx)$ (3.2)
@@
$V^{N}(s,w,\boldsymbol{x})+\frac{1}{2\epsilon}|s-t|^{2}+\frac{1}{2\epsilon}|w-z|^{2}+\frac{1}{2\epsilon}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu)-\Phi(t,z,\mu)$
@@
$=V^{N}(s_{0},w_{0},\boldsymbol{x}_{0})+\frac{1}{2\epsilon}|s_{0}-t_{0}|^{2}+\frac{1}{2\epsilon}|w_{0}-z_{0}|^{2}+\frac{1}{2\epsilon}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})-\Phi(t_{0},z_{0},\mu_{0}).$ (3.3)
@@
Then inequality (3) implies that
@@
$\geq V^{N}(t_{0},w_{0},\boldsymbol{x}_{0})-\Phi^{\epsilon}(t_{0},w_{0},\boldsymbol{x}_{0}).$ (3.5)
@@
$ND_{x^{i}}\Phi^{\epsilon}(s_{0},w,\boldsymbol{x})$ $=\frac{1}{2\epsilon N}D_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu_{0})(x^{i}),\quad\quad\quad\quad\ i=1,\dotso,N,$
$ND_{x^{i}}^{2}\Phi^{\epsilon}(s_{0},w,\boldsymbol{x})$ $=\frac{1}{2\epsilon N}D^{2}_{x\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu_{0})(x^{i}),\quad\quad\quad\quad i=1,\dotso,N,$
@@
$D_{\mu}\Phi(t_{0},z_{0},\mu_{0})(x^{i})$ $=\frac{1}{2\epsilon N}D_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}),\quad\quad\quad\quad\ i=1,\dotso,N,$
$D_{x\mu}^{2}\Phi(t_{0},z_{0},\mu_{0})(x^{i})$ $=\frac{1}{2\epsilon N}D^{2}_{x\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}),\quad\quad\quad\quad i=1,\dotso,N.$
@@
Inequality (3) implies that $t_{0}$ is a maximizer of
@@
$\partial_{x^{i}}\Phi^{\epsilon}(s_{0},w,\boldsymbol{x})$ $=-\frac{1}{2\epsilon N}D_{\mu}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu_{0})(x^{i})$
@@
$\partial^{2}_{x^{i}}\Phi^{\epsilon}(s_{0},w,\boldsymbol{x})=\frac{1}{2\epsilon N}D^{2}_{x\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu_{0})(x^{i}).$
Again due to (3), $\mu_{0}$ is a maximizer of
@@
The last claim is a property of super-convolution; see e.g. [17, Lemma A.5], and hence we finish proving the lemma. ∎
Without loss of generality, we assume that $V^{N}-\Phi^{\epsilon}$ obtains a strict local minimal at $(t_{0},w_{0},\boldsymbol{x}_{0})$. A direct application of Jensen’s lemma ([17, Lemma A.3]) and Lemma 3.1 shows that
$(\partial_{t}\Phi,\partial_{\mu}\Phi,\partial_{x\mu}^{2}\Phi,\Delta_{z}\Phi)(t_{0},z_{0},\mu_{0})$
is in the closure of second order jets of $\Phi^{\epsilon}$ at $(s_{0},w_{0},\boldsymbol{x}_{0})$. We summarize this key result in the following proposition.
@@
and $w\mapsto V^{N}(t_{0},w,\boldsymbol{x}_{0})-\Phi^{\epsilon}(t_{0},w,\boldsymbol{x}_{0})$ is second-order differentiable at $w_{n}$. Moreover, the limit point of
$\left(\partial_{s}\Phi^{\epsilon},ND_{x^{i}}\Phi^{\epsilon},ND^{2}_{x^{i}}\Phi^{\epsilon},\Delta_{w}\Phi^{\epsilon}\right)(s_{0},w_{n},\boldsymbol{x}_{0}),\ \ \ i=1,\dotso,N,$
@@
$\left(\partial_{t}\Phi(t_{0},z_{0},\mu_{0}),\partial_{\mu}\Phi(t_{0},z_{0},\mu_{0})(x^{i}_{0}),\partial_{x\mu}^{2}\Phi(t_{0},z_{0},\mu_{0})(x^{i}_{0}),\Delta_{z}\Phi(t_{0},z_{0},\mu_{0})\right),\ \ \ i=1,\dotso,N.$
@@
$\begin{cases}-\partial_{t}V^{N}(t,w,\boldsymbol{x})=\frac{1}{N}\sum_{i=1}^{N}H^{e}(x^{i},w,ND_{x^{i}}V^{N}(t,w,\boldsymbol{x}),\mu^{\boldsymbol{x}})+\sum_{i=1}^{N}\Delta_{x^{i}}V^{N}(t,w,\boldsymbol{x})\\ \quad\quad\quad\quad\quad\quad\ \ \ \ \ \ +a\sum_{i,j=1}^{N}\operatorname{tr}(D_{x^{i}x^{j}}V^{N})(t,w,\boldsymbol{x}),\\ \quad\ V^{N}(T,w,\boldsymbol{x})=G^{e}(w,\mu^{\boldsymbol{x}}),\end{cases}$
and thus also a viscosity solution. Remember that $V^{N}-\Phi^{\epsilon}$ obtains a local minimal at $(s_{0},w_{0},\boldsymbol{x}_{0})$ in (3), and therefore due to Proposition 3.1
$-\partial_{t}\Phi(t_{0},z_{0},\mu_{0})\geq$ $\frac{1}{N}\sum_{i=1}^{N}H^{e}(x_{0}^{i},w_{0},\partial_{\mu}\Phi(t_{0},z_{0},\mu_{0})(x^{i}_{0}),\mu^{\boldsymbol{x}_{0}})+\sum_{i=1}^{N}\operatorname{tr}{\partial_{x\mu}^{2}\Phi}(t_{0},z_{0},\mu_{0},x_{0}^{i})$
$+a\Delta_{z}\Phi(t_{0},z_{0},\mu_{0}).$
Let us denote $(t_{0},z_{0},\mu_{0})$ by $\theta_{0}$. Comparing the inequality above with (3.2), we define the error term $E(N,\epsilon)$
$E(N,\epsilon):=$ $\sum_{i=1}^{N}\left(H^{e}(x_{0}^{i},z_{0},\partial_{\mu}\Phi(\theta_{0})(x^{i}_{0}),\mu^{\boldsymbol{x}_{0}})-\frac{1}{N}H^{e}(x_{0}^{i},w_{0},\partial_{\mu}\Phi(\theta_{0})(x^{i}_{0}),\mu^{\boldsymbol{x}_{0}})\right)$
$+$ $\int H^{e}(x,z_{0},\partial_{\mu}\Phi(\theta_{0})(x),\mu^{\boldsymbol{x}_{0}})\,\mu_{0}(dx)-\frac{1}{N}\sum_{i=1}^{N}H^{e}(x_{0}^{i},z_{0},\partial_{\mu}\Phi(\theta_{0})(x^{i}_{0}),\mu^{\boldsymbol{x}_{0}})$
$+$ $\int H^{e}(x,z_{0},\partial_{\mu}\Phi(\theta_{0})(x),\mu_{0})\,\mu_{0}(dx)-\int H^{e}(x,z_{0},\partial_{\mu}\Phi(\theta_{0})(x),\mu^{\boldsymbol{x}_{0}})\,\mu_{0}(dx)$
$+$ $\int\operatorname{tr}(D_{x\mu}^{2}\Phi)(\theta_{0},x)\,\mu_{0}(dx)-\sum_{i=1}^{N}\operatorname{tr}{\partial_{x\mu}^{2}\Phi}(\theta_{0},x_{0}^{i}).$ (3.6)
It can be easily checked that (3.2) is satisfies with such $E(N,\epsilon)$. We will provide an upper bound of (3). To this end, let us first estimate $|s_{0}-t_{0}|$, $|z_{0}-w_{0}|$ and $\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})$.
@@
$\frac{1}{2\epsilon}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})-\frac{1}{2\epsilon}\rho_{*}^{2}(\mu^{\boldsymbol{x}^{\prime}},\mu_{0})\leq$ $V^{N}(s_{0},w_{0},\boldsymbol{x}^{\prime})-V^{N}(s_{0},w_{0},\boldsymbol{x}_{0})\leq C(\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu^{\boldsymbol{x}^{\prime}})+\alpha(N))$
$\leq$ $C(\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+2\alpha(N)).$
Hence we obtain an inequality

### v2

@@ 1. Introduction
$\begin{cases}-\partial_{t}v^{N}(t,\boldsymbol{x})=\frac{1}{N}\sum_{i=1}^{N}H(x^{i},ND_{x^{i}}v^{N}(t,\boldsymbol{x}),\mu^{\boldsymbol{x}})+\sum_{i=1}^{N}\Delta_{x^{i}}v^{N}(t,\boldsymbol{x})\\ \quad\quad\quad\quad\quad\quad\ \ +a\sum_{i,j=1}^{N}\operatorname{tr}(D^{2}_{x^{i}x^{j}}v^{N})(t,\boldsymbol{x}),\\ \quad\ v^{N}(T,\boldsymbol{x})=G(\mu^{\boldsymbol{x}})\end{cases}$ (1.2)
@@
$\begin{cases}-\partial_{t}V(t,z,\mu)=\int_{\mathbb{T}^{d}}H^{e}(x,z,D_{\mu}V(t,z,\mu,x),\mu)\,\mu(dx)\\ \ \,\quad\quad\quad\quad\quad\ \ \ \ +\int_{\mathbb{T}^{d}}\operatorname{tr}(D_{x\mu}^{2}V)(t,z,\mu,x)\,\mu(dx)+a\Delta_{z}V(t,z,\mu),\\ \quad\ V(T,z,\mu)=G^{e}(z,\mu),\end{cases}$ (1.3)
@@
is a viscosity supersolution to (1.3) up to some error term $E(\epsilon,N)$, and therefore by comparison for all $(t,z,\boldsymbol{x})\in[0,T]\times\mathbb{T}^{d+dN}$ we have
$V^{N}(t,z,\boldsymbol{x})+E(\epsilon,N)\geq\overline{V}^{N,\epsilon}(t,z,\mu^{\boldsymbol{x}})+E(\epsilon,N)\geq V(t,z,\mu^{\boldsymbol{x}}).$
@@
Then choosing $\epsilon$ properly as a function of $N$, one obtains that $E(\epsilon,N)\approx C\alpha^{1/3}(N)$ where $C$ is a positive constant independent of $N$ and
$\alpha(N)=\begin{cases}N^{-1/2},\quad&\text{if $d=1$},\\ N^{-1/2}\log(N),&\text{if $d=2$},\\ N^{-1/d},&\text{if $d>2$}.\end{cases}$
@@
    $G$ is $k_{*}$-times continuously differentiable and Lipschitz with respect to $C^{-k_{*}}$;
@@
    $H$ is $k_{*}$-times continuously differentiable in all variables, and satisfies the regularity condition, with $C_{H}>0$
    $\left|H(x,p,\mu)-H(x^{\prime},p^{\prime},\mu^{\prime})\right|\leq C_{H}(1+|p|+|p^{\prime}|)(|x-x^{\prime}|+|p-p^{\prime}|+W_{1}(\mu,\mu^{\prime}))$
    for any $x,x^{\prime}\in\mathbb{T}^{d}$, $p,p^{\prime}\in{\mathbb{R}}^{d}$, $\mu,\mu^{\prime}\in\mathcal{P}(\mathbb{T}^{d})$. For each $R>0$, there is a constant $C_{R}$ such that for each $(x,p)\in\mathbb{T}^{d}\times B_{R}$
@@
Under Assumption 1.1, $v^{N}$ converges to $v$ with the rate $\alpha^{1/3}(N)$
$\sup_{(t,\boldsymbol{x})\in[0,T]\times\mathbb{T}^{dN}}\left|v^{N}(t,\boldsymbol{x})-v(t,\mu^{\boldsymbol{x}})\right|\leq C\alpha^{1/3}(N),$
@@
Our result provides the first convergence rate for particle approximations of partially second-order PDEs on the Wasserstein space. Such equations arise in mean-field control problems when the control of common noise is state-independent, as well as in stochastic control with partial observation; see, e.g., [4]. In the former case, the particle approximation justifies the mean-field formulation of controlling a large population of agents; see, e.g., [12, 14] and references therein.
Without the common noise, i.e. setting $a=0$ in (1.1), [19] obtains the convergence rate of $1/\sqrt{N}$ under some convexity assumptions. The argument relies on the fact that super-convolution of semi-concave functions provides desired regularity. While [19] shows better convergence rate, our argument is purely analytic and doesn’t rely on the semi-concavity/convexity of $H$. We would like to mention that the argument of our result makes use of the regularity of $v^{N}$ established in [20] where only the convergence
$\sup_{(t,\boldsymbol{x})\in[0,T]\times\mathbb{T}^{dN}}\left|v^{N}(t,\boldsymbol{x})-v(t,\mu^{\boldsymbol{x}})\right|\to 0$
is provided but not its convergence rate. Moreover, using the same method of the infinite dimensional inf/sup-convolution and the comparison result from [4], one should also be able to show the convergence rate for the same type of equations on ${\mathbb{R}}^{d}$. But to illustrate the main idea of the argument, we choose to work on $\mathbb{T}^{d}$ to avoid the technical issue of non-compactness.
@@ 1.1. Related Literature
Convergence of particle system in mean field control problems were studied in [36, 33, 2, 15] based on viscosity theory, while [27] provided a probabilistic argument. The convergence rate for first-order PDEs on Wasserstein space was obtained in [11, 19, 9]. Assuming the existence of smooth solution to mean-field PDEs, [25] got the optimal convergence rate by a verification argument.
@@ 1.2. Notations
Throughout the paper, we adopt Lions differentiability for functions defined on ${\mathcal{P}}(\mathbb{T}^{d})$; see e.g. [14, Chapter 5].
@@ 2.1. Definition of viscosity solution
Similarly, we define viscosity supersolution. A continuous function
$v:[0,T]\times{\mathcal{P}}(\mathbb{T}^{d})\to{\mathbb{R}}$
is called a viscosity solution if its extension $V$ via (2.1) is a viscosity subsolution and supersolution to (1.3) at the same time.
@@ 2.2. Preliminary results
As the HJB equation (1.2) of particle system is uniformly elliptic, one can show the regularity of solution. The following result is from [20, Lemma 3.1, Theorem 3.2].
@@
$\sup_{(t,\mu)\in{[0,T]\times\mathbb{T}^{d}}}\lVert\frac{\delta\hat{v}^{N}}{\delta\mu}(t,\mu,\cdot)\rVert_{{k_{*}}}\leq\sum_{k=1,\dotso,k_{*}}\sum_{i=1}^{N}\lVert D_{x^{i}}^{k}v^{N}\rVert_{\infty}\leq C.$
@@
By Arzelà–Ascoli theorem, Lemma 2.2 implies there exists a subsequence of $\hat{v}^{N}$ uniformly converging to a function $v$, which is also $\rho_{*}$-Lipschitz. Moreover according to Proposition 2.1, $v$ is also a limiting point of $v^{N}$, and hence is a viscosity solution to (1.1) by a standard argument. Finally, the uniqueness of viscosity solution has been proved in [20] adapting the techniques in [4].
@@
Under Assumption 1.1 for (1.1), any upper semi-continuous subsolution to is smaller than lower semi-continuous supersolution. Then there is a unique viscosity solution $v$ to (1.1) on $\mathcal{P}(\mathbb{T}^{d})$.
@@
Thanks to Sobolev embedding theorem [7, Corollary 9.13], as $k_{*}=\lfloor d/2\rfloor+2$ any $f\in\mathbb{H}^{k_{*}}(\mathbb{T}^{d})$ is Lipschitz with coefficient proportional to $\lVert f\rVert_{{k_{*}}}$. Therefore, it can be easily seen that
@@
$V^{N}(t,z,\boldsymbol{x})\geq V(t,z,\mu^{\boldsymbol{x}})-C\alpha^{1/3}(N),$ (3.1)
@@
$V^{N}(t,z,\boldsymbol{x})\leq V(t,z,\mu^{\boldsymbol{x}})+C\alpha^{1/3}(N)$
@@
$-\partial_{t}\Phi(t_{0},z_{0},\mu_{0})+E(N,\epsilon)\geq$ $\int_{\mathbb{T}^{d}}H^{e}(x,z_{0},D_{\mu}\Phi(t_{0},z_{0},\mu_{0},x),\mu_{0})\,\mu_{0}(dx)$
$+\int_{\mathbb{T}^{d}}\operatorname{tr}(D_{x\mu}^{2}\Phi)(t_{0},z_{0},\mu_{0},x)\,\mu_{0}(dx)+a\Delta_{z}\Phi(t_{0},z_{0},\mu_{0}),$ (3.2)
@@
$V^{N}(s,w,\boldsymbol{x})+\frac{1}{2\epsilon}\left(|s-t|^{2}+|w-z|^{2}+\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu)\right)-\Phi(t,z,\mu)$ (3.3)
@@
$=V^{N}(s_{0},w_{0},\boldsymbol{x}_{0})+\frac{1}{2\epsilon}\left(|s_{0}-t_{0}|^{2}+|w_{0}-z_{0}|^{2}+\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})\right)-\Phi(t_{0},z_{0},\mu_{0}).$
@@
Then inequality (3.3) implies that
@@
$\geq V^{N}(t_{0},w_{0},\boldsymbol{x}_{0})-\Phi^{\epsilon}(s_{0},w_{0},\boldsymbol{x}_{0}).$ (3.5)
@@
$ND_{x^{i}}\Phi^{\epsilon}(s_{0},w,\boldsymbol{x})$ $=\frac{1}{2\epsilon}D_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu_{0})(x^{i}),\quad\quad\quad\quad\hskip 86.0pt\ \,i=1,\dotso,N,$
$ND_{x^{i}}^{2}\Phi^{\epsilon}(s_{0},w,\boldsymbol{x})$ $=\frac{1}{2\epsilon}D^{2}_{x\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu_{0})(x^{i})+\frac{1}{2\epsilon N}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i},x^{i}),\,\,i=1,\dotso,N,$
@@
$D_{\mu}\Phi(t_{0},z_{0},\mu_{0})(x^{i})$ $=\frac{1}{2\epsilon}D_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}),\quad\quad\quad\quad\ i=1,\dotso,N,$
$D_{x\mu}^{2}\Phi(t_{0},z_{0},\mu_{0})(x^{i})$ $=\frac{1}{2\epsilon}D^{2}_{x\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}),\quad\quad\quad\quad i=1,\dotso,N.$
@@
Inequality (3.3) implies that $t_{0}$ is a maximizer of
@@
$D_{x^{i}}\Phi^{\epsilon}(s_{0},w,\boldsymbol{x})$ $=-\frac{1}{2\epsilon N}D_{\mu}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu_{0})(x^{i})$
@@
$D^{2}_{x^{i}}\Phi^{\epsilon}(s_{0},w,\boldsymbol{x})=\frac{1}{2\epsilon N}D^{2}_{x\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}},\mu_{0})(x^{i})+\frac{1}{2\epsilon N^{2}}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i},x^{i}).$
Again due to (3.3), $\mu_{0}$ is a maximizer of
@@
The last claim is a property of super-convolution; see e.g. [18, Lemma A.5], and hence we finish proving the lemma. ∎
Without loss of generality, we assume that $V^{N}-\Phi^{\epsilon}$ obtains a strict local minimal at $(s_{0},w_{0},\boldsymbol{x}_{0})$. Denoting $\theta_{0}=(t_{0},z_{0},\mu_{0})$, a direct application of Jensen’s lemma ([18, Lemma A.3]) and Lemma 3.1 shows that
$\left(\partial_{t}\Phi(\theta_{0}),D_{\mu}\Phi(\theta_{0})(x^{i}_{0}),D_{x\mu}^{2}\Phi(\theta_{0})(x^{i}_{0})+\frac{1}{2\epsilon N}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}_{0},x^{i}_{0}),\Delta_{z}\Phi(\theta_{0})\right).$
approximate derivatives of $\Phi^{\epsilon}$ at $(s_{0},w_{0},\boldsymbol{x}_{0})$. We summarize this key result in the following proposition.
@@
and $w\mapsto V^{N}(t_{0},w,\boldsymbol{x}_{0})-\Phi^{\epsilon}(t_{0},w,\boldsymbol{x}_{0})$ is second-order differentiable at $w_{n}$. Moreover, for all $i=1,\dotso,N$, the limit point of
$\left(\partial_{s}\Phi^{\epsilon},ND_{x^{i}}\Phi^{\epsilon},ND^{2}_{x^{i}}\Phi^{\epsilon},\Delta_{w}\Phi^{\epsilon}\right)(s_{0},w_{n},\boldsymbol{x}_{0}),$
@@
$\left(\partial_{t}\Phi(\theta_{0}),D_{\mu}\Phi(\theta_{0})(x^{i}_{0}),D_{x\mu}^{2}\Phi(\theta_{0})(x^{i}_{0})+\frac{1}{2\epsilon N}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}_{0},x^{i}_{0}),\Delta_{z}\Phi(\theta_{0})\right).$
@@
$\begin{cases}-\partial_{t}V^{N}(t,w,\boldsymbol{x})=\sum_{i=1}^{N}\frac{H^{e}(x^{i},w,ND_{x^{i}}V^{N}(t,w,\boldsymbol{x}),\mu^{\boldsymbol{x}})}{N}+\sum_{i=1}^{N}\Delta_{x^{i}}V^{N}(t,w,\boldsymbol{x})\\ \quad\quad\quad\quad\quad\quad\ \ \ \ \ \ +a\Delta_{w}V^{N}(t,w,\boldsymbol{x}),\\ \quad\ V^{N}(T,w,\boldsymbol{x})=G^{e}(w,\mu^{\boldsymbol{x}}),\end{cases}$
and thus also a viscosity solution. Remember that $\theta_{0}=(t_{0},z_{0},\mu_{0})$ and $V^{N}-\Phi^{\epsilon}$ obtains a local minimal at $(s_{0},w_{0},\boldsymbol{x}_{0})$ in (3). Setting $\epsilon=\alpha(N)$ in Lemma 3.2 below, it can be seen that $s_{0}<T$ for large enough $N$. Therefore due to Proposition 3.1 and the definition of viscosity solution
$-\partial_{t}\Phi(\theta_{0})\geq$ $\frac{1}{N}\sum_{i=1}^{N}H^{e}(x_{0}^{i},w_{0},D_{\mu}\Phi(\theta_{0})(x^{i}_{0}),\mu^{\boldsymbol{x}_{0}})$
$+\sum_{i=1}^{N}\left(\frac{1}{N}\operatorname{tr}{D_{x\mu}^{2}\Phi}(\theta_{0},x_{0}^{i})+\frac{1}{2\epsilon N^{2}}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}_{0},x^{i}_{0})\right)+a\Delta_{z}\Phi(\theta_{0}).$
Comparing the inequality above with (3), we define the error term $E(N,\epsilon)$
$E(N,\epsilon):=$ $\frac{1}{N}\sum_{i=1}^{N}H^{e}(x_{0}^{i},z_{0},D_{\mu}\Phi(\theta_{0})(x^{i}_{0}),\mu^{\boldsymbol{x}_{0}})-H^{e}(x_{0}^{i},w_{0},D_{\mu}\Phi(\theta_{0})(x^{i}_{0}),\mu^{\boldsymbol{x}_{0}})$
$+$ $\int H^{e}(x,z_{0},D_{\mu}\Phi(\theta_{0})(x),\mu^{\boldsymbol{x}_{0}})\,\mu_{0}(dx)-\sum_{i=1}^{N}\frac{H^{e}(x_{0}^{i},z_{0},D_{\mu}\Phi(\theta_{0})(x^{i}_{0}),\mu^{\boldsymbol{x}_{0}})}{N}$
$+$ $\int(H^{e}(x,z_{0},D_{\mu}\Phi(\theta_{0})(x),\mu_{0})-H^{e}(x,z_{0},D_{\mu}\Phi(\theta_{0})(x),\mu^{\boldsymbol{x}_{0}}))\,\mu_{0}(dx)$
$+$ $\int\operatorname{tr}(D_{x\mu}^{2}\Phi)(\theta_{0},x)\,(\mu_{0}-\mu^{\boldsymbol{x}_{0}})(dx)-\sum_{i=1}^{N}\frac{1}{2\epsilon N^{2}}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}_{0},x^{i}_{0}).$ (3.6)
It can be easily checked that (3) is satisfied with such $E(N,\epsilon)$. We will provide an upper bound of (3). To this end, let us first estimate $|s_{0}-t_{0}|$, $|z_{0}-w_{0}|$ and $\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})$.
@@
$\frac{1}{2\epsilon}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})-\frac{1}{2\epsilon}\rho_{*}^{2}(\mu^{\boldsymbol{x}^{\prime}},\mu_{0})\leq V^{N}(s_{0},w_{0},\boldsymbol{x}^{\prime})-V^{N}(s_{0},w_{0},\boldsymbol{x}_{0})$
$\leq C(\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu^{\boldsymbol{x}^{\prime}})+\alpha(N))\leq C(\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+2\alpha(N)).$
Hence we obtain the inequality